eBook Cost Assignment Methods Brody Company makes industrial deaning solvents. Various chemicals, detergent, and water are mixed together and then bottled in 10-gallon provided the following information for last year Raw materials purchases Direct labor Depreciation on factory equipment Depreciation on factory building Depreciation on headquarters building Factory insurance Property taxes: $250,000 140,000 45,000 30,000 60,000 15,000 20,000 18,000 34,000 1,800 150,000 153,000 90,000 127,000 124,000 84,000 102,000 Factory utilities for factory Utilities for sales office Administrative Indirect labor salaries Sales office salaries Beginning balance, raw materials Beginning balance, work in process Beginning balance, finished goods Ending balance, raw materials salaries Check Hy Work 3 more Check My Work uses remaining Next 4) ENG 13/01/2019 1015 PM
media%2F72f%2F72f83bb9-6cd0-4886-895c-9b
eBook Print Item e a cost of goods manufactured statement for last year. Brody Company Statement of Cost of Goods Manufactured For Last Year Total manufacturing cost Calculate the unit product cost. If required, round your answer to the nearest cent 5. Prepare a cost of goods sold statement for last year Brody Company Statement of Cost of Goods Sold For Last Year Check My Work 3 more Check My Work uses remaining
media%2Fb2a%2Fb2a6ab03-65fc-4741-b72e-fc
eBook Print item Statement of Cost of Goods Sold For Last Year 6. Prepare an income statement for last year. Show the percentage of sales that each line item represents. When required, round the percent to four decimal places before converting to a percentage. For example, .88349 would be rounded to.8835 and entered as 88.35. Brody Company Income Statement For Last Year Percent Less: Operating expenses Next Check My Work 8 more Check My Work uses remaining 1016 PM 13/01/2019
Answer #1
1) Direct materials used in production = Beginning balance of raw materials + Purhcases of raw materials - Ending balance of raw materials = 127000+250000-102000 = 275000
2) Total prime cost = Direct materials used + Direct labor cost = 275000+140000 = 415000
3) Total converstion cost = Direct labor cost + Manufacturing ovehead = 140000+(45000+30000+15000+20000+34000+153000) = 437000
4) COST OF GOODS MANUFACTURED
Direct materials used 275000
Direct labor 140000
Manufacturing overhead 297000
Total manufacturing cost 712000
Add: Beginning WIP 124000
Less: Ending WIP 130000
Cost of goods manufactured 706000
Unit product cost = 706000/100000 = 7.06
5) COST OF GOODS SOLD
Beginning finished goods 84000
Add: Cost of goods manufactued 706000
Less: Ending finished goods 79000
Cost of goods sold 711000
6) INCOME STATEMENT
Sales 1200000 100.00%
Cost of goods sold 711000 59.25%
Gross profit 489000 40.75%
Less: Operating expenses:
Depreciation on headquarters building 60000 5.00%
Property taxes-Headquarters 18000 1.50%
Utilities for sales office 1800 0.15%
Administrative salaries 150000 12.50%
Sales office salaries 90000 7.50%
Sales commissions (5% of 120000) 60000 5.00%
Total operating expenses 379800 31.65%
Net operating income 109200 9.10%